# VQ-HPS: Human Pose and Shape Estimation in a Vector-Quantized Latent Space

### v1

@@ 4.3 Training VQ-HPS
VQ-HPS is trained in a supervised manner, given a dataset of RGB images paired with meshes. The canonical mesh predictor is trained solely on the discrete latent representation of meshes. For obtaining the latent representation of the ground truth and for decoding the predicted indices to a full mesh, we use the Mesh-VQ-VAE, which is pre-trained and frozen during the VQ-HPS training.
Mesh-VQ-VAE. The Mesh-VQ-VAE (see Fig. 2) is trained on the AMASS [43] dataset and finetuned on the 3DPW [63] training set. In order to ease the learning of the mesh discrete representation with a limited number of indices, we train the Mesh-VQ-VAE with non-oriented meshes translated to the origin (canonical meshes). The final reconstruction error is 4.7mm. This reconstruction error is an important parameter as it corresponds to the minimal per-vertex error (see Sec. 5.2) we can obtain. Qualitative results of reconstruction on 3DPW are shown in Fig. 3.
Latent canonical meshes. For learning to predict the pose and shape, we only use the discrete representation of the canonical mesh as the training target. The loss is the cross-entropy between the discrete latent representation of the ground truth mesh and the prediction:
$\mathcal{L}_{mesh}=\texttt{Cross-entropy}(\hat{z}_{log},z_{q}),$ (1)
where $\hat{z}_{log}$ are the predicted logits, and $z_{q}$ is the discrete latent representation of the ground truth mesh.
Mesh rotation. We learn to predict the global orientation by computing a mean squared error between the ground truth and predicted rotation matrices:
$\mathcal{L}_{rot}=||\hat{R}-R||_{2},$ (2)
where $\hat{R}$ and $R$ are the predicted and ground truth rotation matrices.
Reprojection. To guide the rotation learning and for better image alignment, we add a reprojection error. It is computed between the 2D projection (using the predicted weak-perspective camera) of the 3D joints extracted from the predicted mesh and the 2D ground truth joints. This loss is computed using the SMPL 24 joints, which can be extracted from the full mesh using a joint regressor $\mathcal{J}_{smpl}$ (see Sec. 3). The reprojection loss is computed as:
@@
In order to improve the final prediction, we propose to add a final rotation predictor network (VQ-HPS-rot). This network takes as input the rotation image features $X_{rot}$, the predicted rotation $\hat{R}$, the predicted camera $\hat{\pi}$, and the predicted skeleton $\hat{p}$ extracted from the predicted mesh $\hat{V}$ using a joints extractor. It outputs a final rotation $\hat{R}_{f}$, which can be applied to the predicted canonical mesh $\hat{V}_{c}$ to predict the final mesh $\hat{V}_{f}$. Fig. 2 overviews this process.
@@ 5 Results
Figure 4: Qualitative results We compare our method with HMR [25], Cliff [36] and FastMetro [7].
@@ 5.1 Datasets
AMASS. The Mesh-VQ-VAE is trained on AMASS [43], which is a large database of human motion in the SMPL [40] format. It contains more than 11000 motions and 300 subjects, which makes it representative of the variety of body poses and shapes.
3DPW. This dataset [63] consists of 60 in-the-wild RGB videos with 3D ground truth for human bodies. We use the pre-defined splits for training, validation, and testing.
@@ 5.2 Metrics
We use several metrics for evaluating the predictions of VQ-HPS. All of them will be expressed in $mm$ for the whole results section.
@@
5.3 Comparison to other methods
@@
HMR [25] 209.34 177.57 89.31
Cliff [36] 223.90 188.82 89.20
FastMetro [7] 176.34 156.99 104.64
VQ-HPS (ours)  [^1] These are results obtained on the valid test set of 3DPW, which is the standard practice [25, 36, 7]. Evaluating on the full 3DPW test set would give a PVE of 184.08 mm, an MPJPE of 157.49 mm, and a PA-MPJPE of 87.12 mm 163.88 139.80 84.92
VQ-HPS-rot (ours)  [^2] These are results obtained on the valid test set of 3DPW, which is the standard practice [25, 36, 7]. Evaluating on the full 3DPW test set would give a PVE of 180.31 mm, an MPJPE of 153.20 mm, and a PA-MPJPE of 87.12 mm 160.68 136.58 84.92
Table 1: Quantitative results on 3DPW. We compare VQ-HPS with SOTA methods using standard metrics.
@@
We train VQ-HPS on the 3DPW training set to see how it performs when trained on limited data. We compare our performance with HMR [25], Cliff [36], and FastMetro [7]. We chose these 3 models for comparison because HMR is the basic architecture for parametric human mesh recovery, Cliff is the state-of-the-art (SOTA) for parametric HPSE, and FastMetro is the SOTA for non-parametric HPSE and the closest method to ours. For this experiment, the backbone for all networks is ResNet-50 [22] pre-trained on ImageNet [52]. We use this backbone because it is the most used in the literature, and its small size seems an appropriate choice for training with limited data. We use the public implementation of FastMetro, and we adapt the implementations of HMR and Cliff provided by [3]. Quantitative results are shown in Tab. 1, and visualizations are available in Fig. 4. In Tab. 1, VQ-HPS denotes our model without the final rotation predictor presented in Sec. 4.3, while VQ-HPS-rot features this improved rotation prediction.
VQ-HPS achieves state-of-the-art results for MPJPE and PA-MPJPE while being very close to the best results in terms of PVE (see Tab. 1). While HMR and Cliff show lower performance in terms of PVE and MPJPE, FastMetro is not state-of-the-art in terms of PA-MPJPE. Our method is the only one showing good performance for all metrics. Visualization of the results confirms that our method performs best. Despite good metrics, FastMetro produces unsmooth results that do not correspond to human shapes. However, we should note that this is probably due to the limited training set, as the results displayed in the original paper looked realistic. This could also be due to the choice of the backbone, since results shown in [7] were obtained with the biggest version of the model, using HRNet-64 [66].
@@ 5.4 Ablation study
Figure 5: Ablation study. Effect of the “3D loss” ablation. We can see that replacing the cross-entropy with a PVE loss produces unnatural poses, showing the interest of our classification-based approach.
We ablate VQ-HPS architecture and training scheme and present the results in Tab. 2.
The ablation “3D loss” replaces the cross-entropy loss with the PVE $\mathcal{L}_{3D}=||V-\hat{V}||_{2}$ where $V$ is the ground truth mesh vertices and $\hat{V}$ the final prediction. Given the huge decrease in performance (see Fig. 5), we conclude that cross-entropy is a good alternative to 3D losses used in all prior works such as [25, 73, 36, 19, 7]. The use of the PVE produces sequences of indices corresponding to anthropomorphic results.
@@
In “Mesh-VQ-VAE", we train Mesh-VQ-VAE only on the AMASS [43] dataset instead of using 3DPW for finetuning as explained in Sec. 4.2. This highlights the importance of this core component in our approach. As Mesh-VQ-VAE is frozen during the VQ-HPS training (see Sec. 4.2), it is crucial to have a good model, and this could be one of the major areas for improvement in our approach.
Method PVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ VQ-HPS 184.08 157.49 87.12 3D loss 233.81 210.41 148.08 No reprojection 193.34 167.00 90.46 Mesh-VQ-VAE 192.32 165.50 91.90
Table 2: Ablation study. We perform several ablations on the VQ-HPS architecture and training process.
Figure 6: Failure cases. VQ-HPS struggles in certain conditions. It sometimes predicts a mesh rotated by 180 degrees relative to the ground truth, especially for symmetric poses. It can also produce wrong predictions when lighting conditions or visibility are bad.
@@ 6 Conclusion
In this work, we proposed Mesh-VQ-VAE, an autoencoder architecture providing a discrete latent representation of registered human meshes. This discrete representation allowed us to tackle the HPSE problem from a classification perspective. It allowed us to avoid the limitations of parametric and non-parametric HPSE methods described in Sec. 1. We also introduced VQ-HPS, a transformer-based model for solving the proposed HPSE classification problem. While being trained using the cross-entropy loss, VQ-HPS outperforms state-of-the-art methods on the 3DPW benchmark, showing the potential of the classification approach for human pose and shape estimation.
Limitations and future works. Despite showing great performance when trained on the 3DPW training set, further experiments would need to be performed with VQ-HPS in order to show its superiority over state-of-the-art methods. We believe that the next step will be to train VQ-HPS on more data. However, a constraint for training our model is the need for 3D mesh ground truth. A potential solution for this problem would be the use of synthetic datasets such as [3, 47]. The evaluation may also be performed on other datasets such as [23, 27].
Failure cases are shown in Fig. 6. We believe that working on videos would solve the problem of occlusions and global orientation. The use of bigger backbones such as [66] pre-trained on pose estimation task would also improve the results and might be sufficient to solve the poor results in the case of bad lighting. Finally, improving the Mesh-VQ-VAE model would probably allow us to avoid unrealistic features such as non-anthropomorphic hands (see the first prediction of Fig. 6).

### v2

@@ 4.2 Training VQ-HPS
VQ-HPS is trained in a supervised manner, given a dataset of RGB images paired with meshes. The canonical mesh predictor is trained solely on the discrete latent representation of meshes. To obtain the latent representation of the ground truth and decode the predicted indices to a full mesh, we use the Mesh-VQ-VAE, which is pre-trained and frozen during the VQ-HPS training. The fact that the Mesh-VQ-VAE is frozen during the training of VQ-HPS is crucial for making realistic predictions in the context of scarce data. Pre-training acts as a regularization, allowing for the reduction of the amount of training data.
Mesh-VQ-VAE. The Mesh-VQ-VAE (see Fig. 2) is trained on the AMASS [29] dataset and finetuned on the 3DPW [30] training set. To ease the learning of the mesh discrete representation with a limited number of indices, we train the Mesh-VQ-VAE with non-oriented meshes translated to the origin (canonical meshes). The final reconstruction error is 4.7 mm. This reconstruction error is an important parameter as it corresponds to the minimal per-vertex error (see Sec. 5.2) we can obtain. Qualitative reconstruction results on 3DPW are shown in Fig. 3.
Latent canonical meshes. For learning to predict the pose and shape, we only use the discrete representation of the canonical mesh as the training target. The loss $\mathcal{L}_{mesh}$ is the cross-entropy between the discrete latent representation of the ground truth mesh and the prediction.
Mesh rotation. We learn to predict the global orientation by computing the mean squared error between the ground truth and predicted rotation matrices.
Reprojection. We add a reprojection error to guide the rotation learning and for better image alignment. It is computed between the 2D projection (using the predicted weak-perspective camera) of the 3D joints extracted from the predicted mesh and the 2D ground truth joints. This loss is computed using the SMPL 24 joints, which can be extracted from the full mesh using a joint regressor $\mathcal{J}_{smpl}$ (see Sec. 3). The reprojection loss is computed as:
@@ 5.1 Datasets
AMASS. The Mesh-VQ-VAE is trained on AMASS [29], a large human motion database in the SMPL [9] format. It contains more than 11000 motions and 300 subjects, which makes it representative of the variety of body poses and shapes.
3DPW. This dataset [30] consists of 60 in-the-wild RGB videos with 3D ground truth for human bodies. We use the pre-defined splits for training, validation, and testing. Note that when training on a mix of datasets (see Sec. 5.4), we do not finetune models on 3DPW to assess generalization. We also evaluate VQ-HPS on 3DPW-OCC, a benchmark proposed in [78] containing videos of 3DPW with occlusions.
EMDB. EMDB [31] contains 81 indoor and outdoor videos with the ground truth SMPL bodies. We use EMDB1, which consists of the 17 most challenging sequences for testing, and train on the rest of the dataset (referred to as EMDB2 in [31]).
COCO. COCO [79] is a dataset of images annotated with 2D keypoints. For training a human mesh predictor, we follow [33, 2] and use pseudo-ground truth meshes. We use the same annotations as [2], with 28k images.
@@ 5.2 Metrics
We use several metrics to evaluate the predictions of VQ-HPS. All of them will be expressed in millimeters (mm) for the whole results section.
@@
5.3 Training on limited data
Table 1: Results on in-the-wild datasets We compare VQ-HPS with SOTA methods using standard metrics on 3DPW trained with 3DPW (1st col.), 3DPW trained with COCO (2nd col.), EMDB trained with EMDB (3rd col.), and EMDB trained with COCO (4th col.). All results are given in mm.
@@ 5.3 Training on limited data
HMR [1] 209.3 110.3 191.3 170.0 177.6 94.1 172.2 149.2 89.3 57.8 96.2 82.4
Cliff [2] 223.9 105.4 163.3 163.2 188.8 89.3 144.7 143.0 89.2 56.8 86.9 81.6
FastMETRO [3] 176.3 107.8 151.0 143.1 157.0 95.8 132.9 123.9 104.6 57.0 95.5 80.2
VQ-HPS (ours) 163.9 102.9 138.5 152.7 139.8 88.0 117.1 131.1 84.9 53.3 77.5 74.5
Figure 4: Qualitative results We compare our method with HMR [1], Cliff [2] and FastMETRO [3] on 3DPW trained on 3DPW (first 3 rows), and EMDB trained on EMDB.
@@
We train VQ-HPS separately on the 3DPW, COCO, and EMDB training sets (see Sec. 5.1) to see how it performs when trained on limited data. We compare our performance with HMR [1], Cliff [2], and FastMETRO [3] trained with the same data. We chose these 3 models for comparison because HMR is the basic architecture for parametric human mesh recovery, Cliff is the state-of-the-art (SOTA) for parametric HPSE, and FastMETRO is the SOTA for non-parametric HPSE and the closest method to ours. For these experiments, the backbone for all networks is ResNet-50 [80] pre-trained on ImageNet [81]. We use the public implementation of FastMETRO and adapt the HMR and Cliff implementations provided by [60]. For all comparisons, we use FastMETRO-S, as this version is the closest to VQ-HPS. When training on COCO, we propose an improved version of VQ-HPS, replacing the MLP of the latent canonical mesh regressor with a Transformer implementing self-attention between the latent mesh features $z_{feat}$. Quantitative results are shown in Tab. 1, and visualizations are available in Fig. 4.
Overall, VQ-HPS outperforms the SOTA methods significantly when training on 3DPW and EMDB (see Tab. 1, col. 1 and 3). Visualization of the results confirms that our method performs best (see Fig. 4), and we propose a more detailed analysis of the error in Appendix D. HMR and Cliff show realistic predictions but are less accurate than VQ-HPS. Despite rather good metrics, FastMETRO produces non-smooth results that do not correspond to human body shapes. This is probably due to the limited training sets, as the results displayed in the original paper looked more realistic. This highlights a clear limitation of non-parametric approaches predicting the 3D coordinates; in comparison, VQ-HPS needs much less data to provide realistic predictions, probably because learning sequences of indices corresponding to anthropomorphic meshes is easier than understanding the 3D structure of the 3D vertices. Furthermore, VQ-HPS benefits from the large-scale pre-training on human motion datasets, acting as a regularization to allow learning with fewer image data labeled with 3D poses.
The fact that the Mesh-VQ-VAE, whose decoder is an important part of VQ-HPS, is pre-trained on AMASS [29] is a great advantage of our approach. Indeed, we can leverage large amounts of body mesh data not paired with images for training. This is particularly interesting because many body motion data can be generated using animation software or generative models. We could finetune the Mesh-VQ-VAE depending on the target application, with uncommon body shapes [82] or extreme poses [83, 84].
FastMETRO slightly takes the lead in global metrics on the EMDB benchmark when training on COCO. The advantage VQ-HPS had when training with scarce data is less important here. Indeed, COCO has as many images as EMDB or 3DPW, but the diversity is much higher than for video datasets, where there exist important correlations between different images [33].
5.4 Training on large-scale datasets
Table 2: Comparison to the SOTA methods. We evaluate VQ-HPS trained on a mix of datasets (see Sec. 5.1) with standard metrics on 3DPW and EMDB and compare them to the state of the art. On the left part, methods use a ResNet-50 backbone. On the right, models use an HRNet backbone, except TokenHMR [76] that uses a ViT [49] backbone and additional data. All results are given in mm.
(a) ResNet-50 backbone
Method PVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
GraphCMR [19] - - 70.2
I2LMeshNet [63] - 93.2 58.6
FastMETRO [3] 129.4 112.6 68.9
HMR [1] - 130.0 81.3
SPIN [39] 116.4 96.9 59.2
PyMAF [42] 110.1 92.8 58.9
ROMP [47] 105.6 89.3 53.5
DSR [48] 105.8 91.7 54.1
PARE [41] 99.7 82.9 52.3
VQ-HPS (ours) 93.6 79.1 50.4
(b) HRNet backbone
3DPW EMDB
Method PVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ PVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
FastMETRO [3] 121.6 109.0 65.7 119.2 108.1 72.7
ROMP [47] 103.1 85.5 54.9 134.9 112.7 75.2
PARE [41] 97.9 82.0 50.9 133.2 113.9 72.2
Virtual Marker [85] 93.8 80.5 48.9 - - -
Cliff [2] 87.6 73.9 46.4 122.9 103.1 68.8
[…]
VQ-HPS (ours) 84.8 71.1 45.2 112.9 99.9 65.2
Following the standard practice [1, 2, 3], we train VQ-HPS on Human3.6M [77], MPI-INF-3DHP [86], COCO [79], and MPII [87]. For this experiment, we use the same version of VQ-HPS as in Sec. 5.3. We evaluate VQ-HPS on 3DPW [30] and EMDB [31] without finetuning on the 3DPW training set. We only compare VQ-HPS with SOTA models using the same backbone and the same datasets for a fair comparison. We take results from the papers or use the provided implementations and checkpoints for other methods. Note that the results on [3, 85] differ from the papers because we do not finetune the models on the 3DPW [30] dataset before testing. Recent methods using a different backbone such as a Vision Transformer (ViT) [49], as well as additional datasets [88, 60, 61, 89] may obtain better results. However, the comparison would not be fair. Results are shown in Sec. 5.4.
VQ-HPS outperforms all other methods on all 3 metrics, being parametric [1, 39, 42, 47, 48, 41, 2] or non-parametric [3, 85, 19, 63]. Note that there is a large gap between the performance of FastMETRO and Virtual Marker in Sec. 5.4 and the reported results in the original papers. This is because we do not perform finetuning on 3DPW. The authors of FastMETRO acknowledge that methods regressing 3D vertices such as [3, 24, 23, 85] perform poorly on data significantly different from the training set [^1] https://github.com/postech-ami/FastMETRO/issues/13. This limitation of non-parametric methods was also described in [24].
Even though TokenHMR’s backbone is more powerful and the method is trained on additional 2D datasets, VQ-HPS still outperforms TokenHMR [76] on the 3DPW and EMDB datasets. When additionally incorporating Bedlam [60] in the training set, TokenHMR takes the lead on the 3DPW dataset as it obtains 84.6, 71.0, and 44.3 mm for the PVE, MPJPE, and PA-MPJPE, respectively, but VQ-HPS remains competitive using much less training data and a less powerful backbone. On the EMDB dataset, TokenHMR using Bedlam for training obtains 109.4, 91.7, and 55.6 mm for the PVE, MPJPE, and PA-MPJPE, respectively.
@@ 5.5 Ablation study
Table 3: Ablation study. We perform several ablations on the VQ-HPS architecture and training process on the 3DPW dataset. All results are given in mm.
Method PVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
VQ-HPS 176.6 152.0 91.8
SMPL 199.8 171.8 99.3
3D loss 220.3 194.9 144.1
No reprojection 183.9 158.4 95.6
Figure 5: Ablation study. Effect of the “3D loss” ablation. We can see that replacing the cross-entropy with a PVE loss produces unnatural poses, showing interest in the classification-based approach.
We ablate VQ-HPS architecture and training scheme and present the results in Tab. 3. We train and test on the 3DPW [30] dataset for these experiments. Note that for faster experiments, we use early stopping with patience of 10 epochs, which makes the training stop much earlier. The first line of Tab. 3 shows the updated VQ-HPS results.
“SMPL" means predicting the SMPL parameters instead of using our proposed discrete latent representation. For obtaining the final 3D prediction, the SMPL model is used instead of the Mesh-VQ-VAE decoder. The performance gap shows that using similar architecture, predicting the discrete latent representation instead of the SMPL parameters yields improved performance.
The ablation “3D loss” replaces the cross-entropy loss with the PVE $\mathcal{L}_{3D}=||V-\hat{V}||_{2}$ where $V$ is the ground truth mesh vertices and $\hat{V}$ the final prediction. Given the huge decrease in performance (see Tab. 3), we conclude that cross-entropy is a good alternative to 3D losses used in all prior works such as [1, 2, 3]. Training VQ-HPS with the PVE produces sequences of indices corresponding to non-anthropomorphic results, which recall results obtained with FastMETRO in Fig. 4 when training on limited data.
@@ 6 Conclusion
In this work, we proposed Mesh-VQ-VAE, an autoencoder architecture providing a discrete latent representation of registered human meshes. This discrete representation allowed us to tackle the HPSE problem from a classification perspective, avoiding the limitations of parametric and non-parametric HPSE methods described in Sec. 1. We also introduced VQ-HPS, a Transformer-based model for solving the proposed HPSE classification problem.
While trained using the cross-entropy loss, VQ-HPS significantly outperforms state-of-the-art methods when trained on scarce data and shows promising performance on large-scale datasets. The classification-based approach exploits the discriminative power of Transformers. It addresses several known problems of non-parametric approaches, such as the plausibility of results and the lack of generalization when training on large-scale datasets without finetuning on the target domain. Comparisons to [3] as well as our ablation study (Sec. 5.5) showed the superiority of the classification-based approach compared to parametric and existing non-parametric approaches when using similar architectures.
Failure cases are discussed in Appendix E. Future works may include the use of a SOTA backbone [49, 90] or additional datasets [89, 60] for achieving better performance. Extensions of the classification-based approach may also be explored for other types of registered meshes, such as human hands or faces. We also believe that multimodal applications involving text and 3D humans [91, 92] would benefit from the Mesh-VQ-VAE representation as it can be considered a language.
Supplementary material
Appendix A Implementation details
In this section, we provide the details of the implementation of Mesh-VQ-VAE and VQ-HPS. More information on these models can be found in the main paper (see Sec. 4.1 and Fig. 2).
[…]
Appendix B Qualitative results
Some qualitative results obtained with VQ-HPS trained on real image datasets and finetuned on the 3DPW training set (see Sec. 5.4, quantitative results in the "With finetuning" part of Sec. 5.4) are shown in Fig. 1.
Figure 1: Qualitative results. We visualize results obtained with VQ-HPS on the 3DPW dataset.
Appendix C Visualization of predictions during training
We visualize validation samples when training on the EMDB dataset (see Sec. 5.3) in Fig. 2. The meshes are smooth from the first epoch and become anthropomorphic in about 5 epochs. As discussed in the paper, we believe this is due to multiple factors. First, the Mesh-VQ-VAE, whose decoder is essential to VQ-HPS, was pre-trained on large-scale human motion datasets and is frozen afterward. This pre-training is probably a regularization that reduces the labeled data needed for learning to solve the HPSE task. Second, VQ-HPS learns a distribution over discrete indices for producing anthropomorphic meshes. Learning a distribution over 54 discrete indices is easier than learning the structure from 6890 3D coordinates.
Figure 2: Visualization of random validation samples during the training process on EMDB. Meshes are smooth from the first epoch and become anthropomorphic in about 5 epochs.
Appendix D Analysis of the error
Figure 3: Distribution of the PVE. We study the distribution of the per-vertex error for all 4 methods on the testing set of EMDB. Black dots represent the mean. The box plots give the 1st and 3rd quartiles, as well as the median value, with a notch for the confidence interval around the median value. The whiskers extend from the box to the farthest data point, lying within 1.5x the interquartile range (IQR) from the box. We also draw the violin plot to visualize the distribution of the error.
We analyze the distribution of the per-vertex error (PVE, see Sec. 5.2) on EMDB1 when models are trained on EMDB2 (see Sec. 5.3, and quantitative results in Tab. 1) in Fig. 3. We can see that VQ-HPS outperforms other methods in many ways. First, its mean, median, 1st, and 3rd quartiles are lower than other methods. The distance between the 1st and 3rd quartiles is much smaller for VQ-HPS than for other methods. Non-parametric methods have a very high concentration of samples with low error, while parametric methods have a more uniform distribution. Another significant advantage of VQ-HPS is its many fewer outliers with high errors.
Appendix E Failure cases
Figure 4: Failure cases. We study the failure cases when training on scarce data and large-scale datasets.
[…]
Appendix F Estimating body shapes
Figure 5: Estimating body shapes. We evaluate VQ-HPS qualitatively on SSP-3D before and after finetuning on a dataset with diverse body shapes.
[…]
Appendix G Visualization of the Mesh-VQ-VAE
Figure 6: Exchanging body parts. We exchange body parts between the two meshes in the latent space after manually identifying the indices responsible for each body part. The color map shows the distance between $\mathcal{M}_{1}$ and the reconstruction.
Figure 7: Exchanging the torso. We exchange the torsos between the two meshes in the latent space.
Figure 8: Interpolation in the latent space of Mesh-VQ-VAE. Note that the latent space encapsulates both pose and shape.
The main paper ( see Sec. 4.1) explains that the Mesh-VQ-VAE is fully convolutional. Hence, the latent space preserves the spatial structure of the mesh. We manually identify the body part associated with each index by visualizing the reconstruction after randomly modifying each. We propose to visualize this property by exchanging body parts between different meshes. Specifically, given two meshes $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, we encode both meshes and decode $\mathcal{M}_{1}$ after replacing the indices of a given body part by $\mathcal{M}_{2}$. Results are shown in Fig. 6.
We can also modify the torso. To visualize it easier, we provide qualitative results for individuals with different body shapes in Fig. 7. Note that this slightly differs from modifying the body shape, as the arms and legs are unchanged.
Finally, in Fig. 8, we show that we can interpolate between meshes $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in the latent space. Specifically, we do a linear interpolation between the corresponding continuous latent representations $z_{1}$ and $z_{2}$, which are then quantized and decoded to obtain intermediate meshes.
[…]